# 2]Huawei Noah’s Ark Lab

### v1

@@ 4.2 Main Results
Since there was no previous work addressing unsupervised lane detection task, to conduct comparison fairly, we directly compare our methods with supervised PointLaneNet [3] as counterpart. The comparison shows that our self-supervised method achieves competitive results compared with supervised methods. On LLAMAS val. set, our method even outperforms its supervised counterpart.
Ablated effects of pre-training on Waymo. To make fair comparison and address the concern of utilizing extra Waymo [54] dataset, we both conduct experiments with and without SLC network from Table 3 to Table 6. The performance of model without SLC network, i.e., naïve lane detector simply trained with pseudo labels from Waymo, denotes the potential effects of pre-training on Waymo dataset. Additionally, the SLC network brings significant improvement to the naïve lane detector, demonstrating the effectiveness of the self-supervised lane detection framework.
TuSimple. For TuSimple benchmark, we report both accuracy and F1 score in Table 3. As demonstrated, the gap between method of strongly supervised and our method is quitely small. Especially, our self-supervised method achieves considerable results at $92.91\%$ F1-measure compared with supervised counterpart, which achieves $95.07\%$ at F1-measure. The comparison result demonstrates the efficiency of our proposed framework.
@@
CurveLanes. Table 6 shows the performance of our method on CurveLanes. As the results show, our LaneCorrect achieves considerable results at $60.39\%$ F1-measure, which is a reasonable result compared with supervised method.
@@ 4.3 Ablation study
Generalization. Inspired by [58], we perform experiments on cross domain tasks. To test the generalization performance of SLC, the student lane detector distilled by our SLC network on CULane domain is evaluated on TuSimple testing set. We also transfer the supervised PointLaneNet model trained on the CULane training set to TuSimple testing set. Table 8 shows that compared with supervised method, LaneCorrect achieves excellent progress in generalization, especially in terms of FPR and FNR. Even compared with other state-of-the-art supervised method, our unsupervised framework achieves comparable results, which proves the generalization ability of our method.
@@
Table 8: Comparison about generalization ability with other supervised methods from CULane training set to TuSimple testing set. “*” represents that the results are from original paper [58]. “-” means that the results are not reported. Our method achieves on par performance with SOTA method at accuracy, but has lower FP and FN rates, resulting a much higher F1 than supervised method.
@@
Table 10: Improvements of LaneCorrect algorithm at each step on Waymo.
@@
Table 11: Ablation studies on hyper-parameter $\lambda_{embed}$.
@@
Extendability. In our main experiments, we choose PointLaneNet as our naïve lane detector because it is a simple and stable lane detection method. Our SLC network is also based on the PointLaneNet. The design of online and target branch enables our SLC network to realize regularization on geometric consistency and contrast learning in feature representation, so as to correct noisy annotations. To clarify the robust and extendability, we also carry out experiments using LaneATT [5] and CLRNet [30] in Table 9.
@@
Table 12: Ablation studies on CNN backbones.
Effects of backbones. To explore the effects of CNN backbones, we conduct the ablation study in Table 12. The largest performance gap, with a difference of 2.08% in F1 score, is observed between ResNet101 [57] and VGG16 [59].
@@ 4.4 Visualization
As Figure 7 demonstrates, with the help of self-supervised trained SLC network, the performance of our unsupervised lane detector is substantially improved. The visualization results are generated on Tusimple dataset. The third column shows scenarios in which unsupervised naïve lane detector predicts incorrectly, and the fourth column shows that with distillation of the self-supervised pre-trained SLC network, our unsupervised lane detector is able to generate correct lane predictions in these cases where naïve lane detector mispredicts.
@@
It is worth noting that, our SLC network significantly improves the performance of model without SLC, which proves the effectiveness of our self-supervised approach.
Generalization. Here we show some qualitative results on cross domain task in Figure 8. The third column presents the visualization results of supervised method, PointLaneNet [3], which is strongly supervised trained on the CULane training set and then directly test on TuSimple test set. To test the generalization performance of LULA, we distill SLC network to a student lane detector on CULane and then directly evaluated on TuSimple test set. The performance is shown in the fourth column.
@@ 5 Conclusions
We have shown the LiDAR view for the lane instance can be utilized as pseudo labels for lane detection task. We have proposed a novel self-supervised training scheme for automatically correcting the noisy pseudo annotation with consistency and contrastive representation learning. To perform lane detetion, we distill the pre-trained correction model to a student lane detector on arbitrary target lane benchmark without touching its ground-truth. Our model achieves comparable performance with existing supervised counterpart (pre-trained on ImageNet [60]) on four major benchmarks and shows superior performance on the domain adaptation setting. We believe self-supervised learning method on lane detection can serve as a strong competitor to the supervised rivals for its scalability and generalizability. The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
6 Data Availability Statement
The datasets generated during and/or analysed during the current study are available in the Waymo [54], TuSimple [55], CULane [27], LLAMAS [56] and CurveLanes [4].

### v2

@@ 4.2 Main Results
Since there was no previous work addressing unsupervised lane detection task, to conduct comparison fairly, we directly compare our methods with supervised PointLaneNet Chen et al. (2019b) as counterpart. The comparison shows that our self-supervised method achieves competitive results compared with supervised methods. On LLAMAS val. set, our method even outperforms its supervised counterpart.
Ablated effects of pre-training on Waymo. To make fair comparison and address the concern of utilizing extra Waymo Sun et al. (2020) dataset, we both conduct experiments with and without SLC network from Table 3 to Table 7. The performance of model without SLC network, i.e., naïve lane detector simply trained with pseudo labels from Waymo, denotes the potential effects of pre-training on Waymo dataset. Additionally, the SLC network brings significant improvement to the naïve lane detector, demonstrating the effectiveness of the self-supervised lane detection framework.
TuSimple. For TuSimple benchmark, we report both accuracy and F1 score in Table 3. As demonstrated, the gap between method of strongly supervised and our method is quitely small. Especially, our self-supervised method achieves considerable results at $92.91\%$ F1-mea-  sure compared with supervised counterpart, which achie-  ves $95.07\%$ at F1-measure. The comparison result demonstrates the efficiency of our proposed framework.
We also compare our method with other comparative learning-based approaches in Table 4. The first column indicates the proportion of labeled data used for training, while the remaining data is treated as unlabeled for self or semi-supervised learning. Notably, our method is capable of handling scenarios when there is no (0%) human-labeled data included. Additionally, the efficiency of our method is more pronounced across all situations.
@@
CurveLanes. Table 7 shows the performance of our method on CurveLanes. As the results show, our LaneCorrect achieves considerable results at $60.39\%$ F1-measure, which is a reasonable result compared with supervised method. The visualization results are shown in Fig. 5.
@@ 4.3 Ablation study
Generalization. Inspired by Qu et al. (2021), we perform experiments on cross domain tasks. To test the generalization performance of SLC, the student lane detector distilled by our SLC network on CULane domain is evaluated on TuSimple testing set. We also transfer the supervised PointLaneNet model trained on the CULane training set to TuSimple testing set. Table 9 shows that compared with supervised method, LaneCorrect achieves excellent progress in generalization, especially in terms of FPR and FNR. Even compared with other state-of-the-art supervised method, our unsupervised framework achieves comparable results, which proves the generalization ability of our method.
Table 9: Comparison about generalization ability with other supervised methods from CULane training set to TuSimple testing set. “*” represents that the results are from original paper Qu et al. (2021). “-” means that the results are not reported. Our method achieves on par performance with state-of-the-art method at accuracy, but has lower FP and FN rates, resulting a much higher F1 than supervised method.
@@
Table 11: Improvements of LaneCorrect algorithm at each step on Waymo.
@@
Table 12: Ablation studies on hyper-parameter $\lambda_{embed}$.
@@
Extendability. In our main experiments, we choose PointLaneNet as our naïve lane detector because it is a simple and stable lane detection method. Our SLC network is also based on the PointLaneNet. The design of online and target branch enables our SLC network to realize regularization on geometric consistency and contrast learning in feature representation, so as to correct noisy annotations. To clarify the robust and extendability, we also carry out experiments using LaneATT  Tabelini et al. (2021b) and CLRNet Zheng et al. (2022) in Table 10.
@@
Effects of backbones. To explore the effects of CNN backbones, we conduct the ablation study in Table 13. The largest performance gap, with a difference of 2.08% in F1 score, is observed between ResNet101 He et al. (2016) and VGG16 Simonyan and Zisserman (2014).
@@ 4.4 Visualization
As Figure 7 demonstrates, with the help of self-supervised trained SLC network, the performance of our unsupervised lane detector is substantially improved. The visualization results are generated on Tusimple da-  taset. The third column shows scenarios in which unsupervised naïve lane detector predicts incorrectly, and the fourth column shows that with distillation of the self-supervised pre-trained SLC network, our unsupervised lane detector is able to generate correct lane predictions in these cases where naïve lane detector mispredicts.
@@
It is worth noting that, our SLC network significantly improves the performance of the variant without SLC, which proves the effectiveness of our self-supervised approach.
Generalization. Here we show some qualitative results on cross domain task in Figure 8. The third column presents the visualization results of supervised me-  thod, PointLaneNet Chen et al. (2019b), which is strongly supervised trained on the CULane training set and then directly test on TuSimple test set. To test the generalization performance of LULA, we distill SLC network to a student lane detector on CULane and then directly evaluated on TuSimple test set. The performance is shown in the fourth column.
@@ 5 Conclusions
We have shown the LiDAR view for the lane instance can be utilized as pseudo labels for lane detection task. We have proposed a novel self-supervised training scheme for automatically correcting the noisy pseudo annotation with consistency and contrastive representation learning. To perform lane detetion, we distill the pre-trained correction model to a student lane detector on arbitrary target lane benchmark without touching its ground-truth. Our model achieves comparable performance with existing supervised counterpart (pre-trained on ImageNet Deng et al. (2009)) on four major benchmarks and shows superior performance on the domain adaptation setting. We believe self-supervised learning method on lane detection can serve as a strong competitor to the supervised rivals for its scalability and generalizability. The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
Data Availability Statement
The datasets generated during and/or analysed during the current study are available in the Waymo Sun et al. (2020), TuSimple TuSimple (2019), CULane Pan et al. (2018b), LLAMAS Behrendt and Soussan (2019) and CurveLanes Xu et al. (2020).
Acknowledgments
This work was supported in part by National Natural Science Foundation of China (Grant No. 62376060).